### v1

Abstract
In blind quantum computing, a user with a simple client device can perform a quantum computation on a remote quantum server such that the server cannot gain knowledge about the computation. Here, we numerically investigate hardware requirements for verifiable blind quantum computing using an ion trap as server and a distant measurement-only client. While the client has no direct access to quantum-computing resources, it can remotely execute quantum programs on the server by measuring photons emitted by the trapped ion. We introduce a numerical model for trapped-ion quantum devices in NetSquid, a discrete-event simulator for quantum networks. Using this, we determine the minimal hardware requirements on a per-parameter basis to perform the verifiable blind quantum computing protocol. We benchmark these for a five-qubit linear graph state, with which any single-qubit rotation can be performed, where client and server are separated by 50 km. Current state-of-the-art ion traps satisfy the minimal requirements on a per-parameter basis, but all current imperfections combined make it impossible to perform the blind computation securely over 50 km using existing technology. Using a genetic algorithm, we determine the set of hardware parameters that minimises the total improvements required, finding directions along which to improve hardware to reach our threshold error probability that would enable experimental demonstration. In this way, we lay a path for the near-term experimental progress required to realise the implementation of verifiable blind quantum computing over a 50 km distance.
I Introduction
Quantum computers may outperform classical computers in a variety of tasks [1, 2, 3], but these are so far inaccessible; despite progress in realising these devices across a variety of physical platforms [4, 5, 6, 7], building and running quantum computers is associated with a large financial cost [8, 9, 10]. Classical cloud-based access to quantum servers eliminates the need for users to own large and expensive devices themselves [11], but is unsuitable for use cases involving sensitive data, in which a user requires access to the computational power of the quantum server without revealing the input data, the computation or the output to the owner of the server [12, 13, 14].
Blind quantum computing (BQC) is a technique where a client can perform quantum algorithms at a remote quantum server without the input, the computation, or its outcome being revealed (apart from an upper bound on the size of the computation) [15]. Preferably, the client is realised as cheaply as possible, to help make quantum computing more widely available. Two options for realising such client are for it to have the ability to either send single photons, or measure them [16]. In this work, we assume quantum capabilities of the client to be limited to making measurements.
The initial BQC protocol, based on measurement-based quantum computation [17, 18], was later expanded to include verification to test for correctness [19]. In such verifiable blind quantum computing (VBQC) protocols, the client can abort if the output of certain tests are not as expected.
In realistic near-term noisy quantum devices, imperfections are inevitable. In verifying tests, imperfections and noise in the system can be mistaken for malicious behaviour of the server, resulting in a computation that will be aborted constantly, and the client gaining little information about the operations of the server. To combat this, a noise-robust verified BQC (rVBQC) protocol was introduced [20]. This protocol tolerates imperfections from noise or malicious behaviour provided that the server does not fail more than 25% of verifying tests employed by the client (for the type of computation we study, see Section III for details). The robustness to noise is realised by repeating the computation multiple times and performing classical error correction (majority voting) on the results.
Already, there have been proof-of-principle demonstrations of BQC [21] and rVBQC [22] in laboratory settings. For real-world practicality, however, the client needs to be able to be spatially remote from the server, which will require improvements to existing quantum computing and communication hardware.
In this paper, we determine the requirements for performing rVBQC at a metropolitan scale (i.e., the scale of a large city) of 50 km using a trapped-ion based server and a measurement-only client as depicted in Figure 1, using the 25% error tolerance as a threshold. In measurement-based quantum computing, any single-qubit gate can be performed using a five-qubit linear graph state (i.e., five qubits, each in a superposition state, with controlled-$Z$ operations between them) [23]. We use this five-qubit graph state to benchmark the performance of the protocol. We focus on rVBQC because this is feasibly achievable near-term.
We investigate this implementation of the rVQBC protocol numerically using NetSquid, a discrete event simulator for quantum networks [24]. To this end, we introduce a framework for the modelling of trapped-ion quantum servers, including a NetSquid library [25], along with a model of a measurement-only client. The simulation is hardware motivated and takes a set of hardware parameters as input. Using this, we both:
  1. 1.
    Identify the per-parameter minimal requirements for hardware to allow for rVBQC. In each case, we assume perfect performance of all other parameters apart from fibre attenuation. This gives us a strict lower-bound for each parameter, which is compared to state-of-the-art performance. We show that current state-of-the-art ion traps satisfy absolute minimal requirements on a per-parameter basis, but all current imperfections combined make it impossible to perform rVBQC securely over 50 km using existing technology. These results are summarised in Figure 4;
  2. 2.
    Identify the set of hardware parameters that minimises the total improvement needed over current state-of-the-art parameters to allow for a successful implementation of rVBQC. This reveals which parameters need the most improvement and how far we need to improve them to enable a metropolitan-scale application of rVBQC. We do this by combining our requirement on the error probability as well as the cost ([26, 27]) of a set of hardware parameters into a single-objective minimisation problem. This is fed into a genetic algorithm [28] that minimises the total improvement needed over state-of-the-art performance. These results show there is substantial work left to be done in hardware improvements, and are summarised in Figure 5.
We organise this paper as follows. In Section II, we provide details of the physical setup that we simulate, including physical parameters. Then in Section III, we analyse the two primary results of our numerical analysis discussed briefly above. Section IV details the trapped-ion model, the client’s measurement apparatus, as well as the genetic algorithm and cost function that define our optimisation procedure. Finally, in Section V we discuss possible directions for future work.
II Setup
We simulate a two-party rVBQC setup with a trapped-ion quantum server and a measurement-only client using NetSquid. We investigate the protocol at a metropolitan scale, in which the server and client are separated by 50 km of optical fibre. An overview of the setup is provided in Figure 1. The protocol used here, as in [22], assumes a variation wherein the client uses measurements to perform remote state preparation (RSP) on the server. In RSP, a sender measures part of an entangled state and communicates a classical correction to prepare a target state at a receiver. A formal description of the general protocol can be found in A.
Below, we will outline how normally rVBQC includes both computation and test rounds, and why our analysis only focuses on test rounds (Section II.1). In Sections II.2 through II.5 we describe the steps of such a test round. A visualisation of these steps is given in Figure 2. After this, we introduce the parameter sets that are used as input to the simulation (Section II.6) and the metrics on which we base our analysis (Section II.7). This provides necessary context to understand the results as presented in Section III.
ServerClientPBSQQHSPDSPDCBClassical channelQuantum channel
Figure 1: Setup overview of a trapped-ion quantum server (depicted with 5-qubit linear graph state), connected to a client that measures photons coming in through the quantum channel using half (H) and quarter (Q) waveplates to rotate the qubit state, and measures the qubit using a polarising beam splitter (PBS) and single photon detectors (SPD). The classical channel is used by the client to coordinate with the server. The classical box (CB) is a classical unit that the client uses to choose graph colouring and measurement angles, handle classical communication and perform checks on the results send by the server. The photonic qubits sent out by the server are sent over the quantum channel. The client setup is depicted with more detail to show how this is modelled in the simulation, the exact setup of the server is not shown (i.e., it is kept more general).
II.1 Computation and test rounds
In a normal run of the protocol, one picks a total number of rounds $N$ that are separated into computation rounds and test rounds, as suggested in [20]. The output of the total computation is taken as the majority output of the computation rounds.
For a test round we make use of trap qubits that are prepared in a superposition basis and are decoupled from the rest of the graph by having them surrounded by dummy qubits. Dummy qubits are prepared in the standard basis. The client asks the server to measure the traps in the basis they are prepared in and is therefore aware of what the outcome should be. A test round fails when any of the measurement outcomes of the traps do not equal the expected outcome. A test can fail due to noise or due to malicious behaviour by the server. In the computation considered here, the error tolerance for test rounds is 25%, as is explained in Section IV.3.
In our analysis, we focus only on test rounds as the computation rounds do not provide additional information about the verifiability of the setup. From the test rounds, we extract the error probability as the fraction of failed rounds. With the resource we consider, a five-qubit linear graph (i.e., five qubits connected through $CZ$ gates)— the test rounds can be accompanied with computation rounds in which any single-qubit gate is performed [29]. Thus, our results are universal for single-qubit gates.
We will outline how the test rounds are performed by going through the steps in Sections II.2 through II.5, Figure 2 provides an overview of these steps.
[…]
Figure 2: Visualisation of a test round. Depicts tasks (blocks) performed by the client (left) and server (right) along with communication between them (arrows). Blue blocks and arrows represent classical tasks and communication, orange blocks and arrows represent quantum tasks and communication. Grey blocks add comments and green blocks highlight the measurement outcomes. Depicted in chronological order from top to bottom, where the boxed areas (black outline) are repeated five times, assuming they are successful, as these steps are performed for all qubits in the graph.
II.2 Remote state preparation
The client controls the state of the qubits at the server by performing remote state preparation (RSP). In this work, we assume the quantum capabilities of the client to be limited to making measurements. Using entangled qubit pairs, the client can project a qubit located in the server onto a chosen basis by measuring an entangled qubit emitted (and sent to the client). In this way, the state of the qubits is known to (and determined by) the client, but it is unknown to the server. This is the source of the blindness of this protocol.
In each round, the client randomly chooses the sets of qubits that will become the trap qubits and that will become the dummy qubits, such that all traps are surrounded by dummies. For a five-node graph, one can have either dummy-trap-dummy-trap-dummy or trap-dummy-trap-dummy-trap (other combinations are sub-optimal for trap insertion). The dummy qubits need to be prepared in the standard basis $H/V$ (i.e., on the north or south pole of the Bloch sphere) and the trap qubits need to be prepared in a superposition basis |±$\ket{\pm_{\theta_{v}}}=(\ket{H}\pm e^{i\theta_{v}}\ket{V})/\sqrt{2}$ (i.e., on the equator of the Bloch sphere) where $\theta_{v}\in\{k\pi/4\}_{0\leq k\leq 7}$, with $k$ an integer such that the angle $\theta_{v}$ is randomly and independently chosen for each qubit $v$.
To start off RSP, the client communicates a description of the graph it wants to prepare to the server, this graph is the same for all rounds. This description includes the nodes in the graph $I$ and the edges of the graph $G$ (describing which nodes are connected through $CZ$ gates). Here, we consider a five-qubit linear graph ($I=\{0,1,2,3,4\};G=\{(0,1),(1,2),(2,3),(3,4)\}$).
After receiving the description, the server emits a polarisation-encoded photonic qubit that is entangled with an ion as $|\psi\rangle=(\ket{0H}+\ket{1V})/\sqrt{2}$ and this photonic qubit is sent to the client over 50 km of optical fibre. The client then measures this photon in the standard or superposition basis, depending on whether it is a trap or dummy. The client thereby remotely prepares five qubits on the server that are in the state |$\ket{m_{v}}$ or |$\ket{+_{\theta_{v}^{\prime}}}=(\ket{0}+e^{i\theta_{v}^{\prime}}\ket{1})/\sqrt{2}$, where $\theta_{v}^{\prime}=\theta_{v}+m_{v}\pi$, with $m_{v}$ the outcome of the client’s measurement. Note that the server does not know the basis in which the client measures, and therefore it is unaware of the state of the qubits in its memory. Because the server does not know the state of the qubits, it also cannot tell the difference between a computation round (which involves only |±$\ket{\pm_{\theta_{v}}}$-qubits) and a test round (which includes dummy qubits) without performing a (malicious) intermediate measurement that has a chance of disturbing the quantum state.
The client performs the superposition basis measurement by rotating the polarisation state of the incoming ion-entangled-photonic qubit followed by measuring the qubit in the standard basis. Rotating the polarisation state of the photonic qubit can be done by optical elements such as waveplates or electro-optic modulators. A measurement in the standard basis can be performed by a polarising beam splitter (PBS) followed by single photon detectors (SPD). For a more detailed description of a possible physical realisation of the client, see [22]. The simulated rotation setup consists of one half waveplate, followed by two quarter waveplates based on [30].
II.3 Cutoff time
In the RSP process, some qubits will have been prepared and will be sitting in the memory, waiting for the preparation of the remaining qubits. Depending on the efficiency of the server and due to loss from the fibre, remotely preparing a qubit might take many tries, meaning that the qubits may have to reside in the memory for a long time. The longer the qubits are in memory, the more they decohere (details in Section IV.1). For a lower error probability, we want to limit the time qubits spend in memory, for which we include a cutoff time. This cutoff ensures that no qubit resides in the server memory for longer than half a coherence time during the RSP phase, thereby limiting the decoherence.
Choosing a good cutoff time is a trade-off between rate and fidelity: a shorter cutoff time gives higher quality qubits as they have suffered from less decoherence, yet more qubits will be discarded, leading to a lower rate. However, since this work only targets an error probability (which depends on the fidelity of the remotely prepared qubits) and does not consider rate, no optimisation can be performed on this aspect. See e.g. [31] for examples of optimising this. Instead, we choose a fixed cutoff of half the coherence time.
When any of the qubits in memory have been at the server for longer than the cutoff time, they are discarded and the client is notified. When the server has prepared five qubits, no more qubits will be discarded (i.e., no more cutoff after RSP).
II.4 Graph formation
The client instructs the server to arrange the qubits in the linear graph: edges should be made between qubits $(0,1),(1,2),(2,3)$ and $(3,4)$. The server does this by applying controlled-$Z$ ($CZ$) operations between successive qubits to form a chain-like graph.
We assume that the ion qubits can be individually addressed for rotations $R_{x},R_{y},R_{z}$. This ability is hardware-dependent (it is possible in, for example, [32, 33, 34]). Other options besides direct-addressing include global addressing of ions either by ‘hiding’ other ions (as in [35]), ‘shuttling’ ions to different physical regions (as in [36, 37, 38]) or adding extra rotations (as in [39]), as long as rotations around at least one axis can be individually addressed.
We assume the bichromatic gate $R_{XX}(\theta)=\exp(-i\theta/2\sum_{i<j}\sigma_{x}^{(i)}\sigma_{x}^{(j)})$, also known as the Mølmer-Sørensen gate [40, 41] to be the native entangling operation. We use Qiskit [42] to find how to construct a $CZ$ gate using the operations available for the ion trap and we show this in Figure 3.
In principle, the server does not need to wait for all qubits to be prepared before it starts applying the CZ operations, but it is done so in the simulation for simplicity. The time it takes to prepare the graph state (around 150 $\mu$s) is short compared to the expected time it takes for the qubits to be remotely prepared (in the order of tens of milliseconds  [^1] Calculated as time it takes to perform one attempt (photon round-trip time: 2x50 km / (c/n) $\approx$ 0.50 ms, with c the speed of light and n = 1.4682 the index of refection of telecommunication optical fibre) divided by the average probability of photon arriving at client (server efficiency times probability of photon transmitting through the fibre: 0.1325 $\times$ (50km $\times$ 0.2 dB/km)=0.1. Giving an average probability of photon arriving of 0.01325) giving 0.50/0.01325=37.8 ms.).
Figure 3: Controlled-$Z$ gate compiled using Qiskit for assumed native gateset of the trapped-ion quantum server. The server is assumed to be able to individually address the ions for rotations ($R_{x},R_{y},R_{z}$), and to be able to perform the bichromatic gate ($R_{XX}$, also known as Mølmer-Sørensen gate) as native entangling gate.
II.5 Server measurements and output
Once the graph state formation is complete, the client instructs the server to measure in ±$\pm_{\delta_{v}}$-bases defined by
$\delta_{v}=\theta_{v}+m_{v}\pi+r_{v}\pi,\text{ for trap qubits},$ (1)
$\delta_{v}\in\{k\pi/4\}_{0\leq k\leq 7},\text{ for dummy qubits}.$ (2)
Here, $\theta_{v}$ defines the basis used by the client to perform RSP and $m_{v}$ is the outcome of the RSP measurement of the photonic qubit entangled with ion qubit $v$. $r_{v}$ is a random bit generated by the client to ensure the measurement outcomes appear completely random. The angles at which the dummy qubits are measured is irrelevant to the client; they just needs to be random, such that the server will not be able to detect the qubit as a dummy (which would reveal that the client is executing a test round).
The server communicates the outcome of measuring qubit $v$ to the client as $b_{v}$. For all trap qubits, the client checks if $b_{v}=r_{v}$. The test round succeeds if this is true for all traps. If any of the traps do not satisfy $b_{v}=r_{v}$, the round fails. The round failing or succeeding is the output for each of the test rounds.
II.6 Parameter sets
The steps from the previous subsections are simulated in NetSquid to find the error probability (i.e., the fraction of failed test rounds) for a given setup. The simulation takes a set of hardware parameters as input. We divide these hardware parameters into two sets: one set that we do not vary during optimisation (parameter set 1), and one set that we do vary during optimisation (parameter set 2). The first set contains parameters that are inherent to the setup (fibre length), parameters that are established optical components with little room for improvement (fibre loss, crosstalk in beamsplitters, waveplate error probability, detector dark-count rate) and the duration of operations at the ion trap. The last is not optimised over as these timescales are already very small compared to the time required to perform RSP (which is in the order of 20 ms per qubit, two orders of magnitude larger than the timescale of performing gates and readout). This first set is presented in Table 1.
Parameter Value
Channel length
50 km
Photon loss probability in fibre
0.2 dB/km
Waveplate error probability (fast axis tilt/retardation deviation)
0.001
Dark count probability photondetectors
0.03%
Crosstalk in polarising beam splitter
0.0001
Qubit rotation duration
12 $\mu$s [44]
Entangling gate duration
107 $\mu$s [44]
Ion initialisation duration
300 ns [45]
Photon emission duration
300 ns [45]
Ion qubit readout duration
100 $\mu$s
Table 1: Parameter set 1: Hardware parameters used as simulation input that are not varied over.
The second set of parameters are the optimisation parameters. These are the parameters we vary over in the minimisation methods explained in Section IV.3. In Table 2 we list these parameters as well as what we consider the ‘baseline’. The baseline is current state of the art in long-distance trapped-ion experiments and is used in the minimisation methods to find the difference between what is currently achievable and what is needed to run the protocol at metropolitan distances, as will be explained in Section IV.3. We choose this baseline instead of values from current record experiments (such as the long coherence time of [46], the high gate fidelities of [34, 32] or the high entanglement fidelity of [47]) because we believe it is more realistic for long-distance experiments. We are actively looking ahead to what we need to achieve rVBQC in a metropolitan scale (i.e., over a 50 km distance), instead of focusing on in-lab experiments, as rVBQC is already feasible in a lab setting [22].
Parameter Baseline value
Server efficiency
(= emit $\times$
freq. conversion)
0.1325
(= 0.53[48] $\times$
0.25[49])
Single-qubit gate fidelity
0.99 [50]
Entangling gate fidelity
0.95 [50]
Emission fidelity
0.974 [51]
Coherence time [ms]
62 [44]
Table 2: Parameter set 2: Baseline of long-distance trapped-ion experiments forming state of the art. We vary over these parameters to find the absolute minimal hardware requirements and minimal improvements as outlined in Section IV.3. Server efficiency here refers to the total efficiency of preparing an ion-qubit, emitting an ion-entangled photon, coupling to the fibre (combined giving the number for ‘emit’ in the table) and converting its frequency to the telecom C band at 1550 nm (‘freq. conversion’ in the table). The emission fidelity refers to the fidelity of the ion-photon entangled pair when the photon is emitted.
Figure 4: Absolute minimum a given parameter can be while still staying below the required maximum error probability of 25% assuming that the only other imperfection in the system is photon loss in fibre (orange). These minima are compared to the state of the art as given in Table 2 (blue/purple, hatched). The dashed black line indicates a change in scaling between the fidelities and the coherence time. Below the solid black line there is a comparison between the overall error with all imperfections in the state of the art (yellow) and the threshold error probability of 25% (green hatched); despite each parameter’s baseline value meeting the minimal requirements, together they are insufficient to run the rVBQC protocol.
II.7 Error tolerance and requirements
Here, we justify the threshold and explain the two different optimisation procedures we use to come to two different sets of minimised parameters.
In abstract cryptography (i.e., the theory behind the security proof of the rVBQC protocol [20]), security is defined as the indistinguishability between the real-world implementation of the protocol and the ideal, noiseless resource [52]. When the security of a protocol holds over sequential or parallel repetitions with other protocols, it is said to be composably secure. The rVBQC protocol we study (A) is $\epsilon$-composably-secure (as defined in [20]) with $\epsilon$ exponentially small in the number of rounds if the fraction of failed test rounds over total number of test rounds is bounded by 25% [20]. Details for this are provided Section IV.3.
We can use the error tolerance to find two sets of requirements: absolute minimal hardware requirements and minimal improvements. To find the absolute minimal hardware requirements, we start by setting all hardware parameters to perfect except for photon loss in fibre. That is, for parameter set 1, we set the waveplate errors, darkcount probability and beam splitter crosstalk to 0, the loss in fibre is kept as presented in Table 1 at 0.2 dB/km. For parameter set 2, we set the server efficiency and all fidelities to 1 and we remove the memory decoherence noise model to simulate effectively infinite coherence time. Then one parameter from set 2 is made progressively worse, until the simulated error probability rises above 25%. The last value of the parameter before this happens is taken as the absolute minimal requirement.
For the minimal improvements, we aim to find the least costly improvements needed over the baseline parameters to get the error probability below 25%. The cost of a hardware parameter is given in terms of an improvement factor. This improvement factor quantifies the difference between a given value and the baseline, being 1 for no improvement, and tending to infinity for a parameter tending to perfect (e.g., infinite coherence time, fidelity of 1). More information on the cost function and the improvement factor can be found in Section IV.3. The minimal improvements are found by combining the error tolerance with minimising the improvement factors into a single objective optimisation problem and solving it using a genetic algorithm.
III Results and discussion
Parameter Baseline Minimally improved value
Server efficiency 0.133 0.393
Single-qubit gate fidelity 0.99 0.997
Entangling gate fidelity 0.95 0.988
Emission fidelity 0.974 0.982
Coherence time [ms] 62 124
Figure 5: (a): Minimally improved parameter set required to perform a 5-qubit linear-graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km with a 25% error probability threshold. This set minimises the cost function (11), meaning it is closest to the state-of-the-art baseline of those that meet the requirement. A visual representation of these parameters in terms of an improvement factor is given in the bottom figure. (b): Directions along which hardware must be improved to perform a 5-qubit linear graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km. The further away the line is from the centre towards a given parameter, the larger the improvement that parameter requires. The improvements are given in terms of an improvement factor k, which tends to infinity as a parameter tends to its perfect value and is 1 for no improvement. More information on this can be found in Section IV.3.
We now find requirements for enabling rVBQC with a measurement-only client and a trapped-ion server separated by 50 km of optical fibre, for five-qubit linear graphs. The data generated for these results is available at [53], the code used to generate the data and plots is available at [54].
The absolute minimal requirements found, compared to the baseline of Table 2, are shown in Figure 4. More information on how the values of Figure 4 are extracted from the simulation outcomes and on how their confidence intervals are determined is given in Section IV.3.
From Figure 4 we see that state-of-the-art ion traps satisfy absolute minimal requirements on a per-parameter basis as all blue bars, representing the baseline, are larger than the absolute minimal requirements in orange. However, in Figure 4 we also show the comparison between the full set of baseline parameters as ‘overall error’ in red (i.e., all imperfections combined), with the threshold in green. This shows us that all current imperfections combined make it impossible to perform the blind computation securely over 50 km using existing technology, as all baseline parameters combined lead to an error probability of 50%, which is twice the requirement of 25%.
Note that there is no absolute minimal requirement on the server efficiency, as there is no threshold for rate, only for error probability. If all other parameters are perfect, any server inefficiency combined with fibre attenuation will still lead to a 0% error probability, as the storage of qubits is perfect. (This is not the case for the minimal improvements set, as other parameters will be imperfect. Having imperfect memory efficiency then leads to qubits residing in imperfect memory for longer, thereby suffering more decoherence.) It will only lead to lower rate, not considered here. The difference between the minimal requirement and the state of the art appears particularly large for the coherence time, however, we note that the comparison is skewed as all other parameters are on a scale of zero to one, which is not the case for coherence time. Because of this we chose to separate the coherence time from other parameters with a dashed line, to indicate a different scaling.
We find the minimal improvements, i.e., the set that is closest to the baseline in Table 2 satisfying the threshold, by setting up and minimising a cost function as defined in (11). More details on the cost function and how the minimisation is performed can be found in Section IV.3. The set of hardware parameters that minimises this cost function are given in Figure 5 (a) and visualised in Figure 5 (b). The plot in Figure 5 (b) shows the improvement factor for each parameter in Figure 5 (a). The improvement factor tends to infinity as a parameter tends to its perfect value and is 1 when no improvement is made compared to the baseline. The further away the line is from the centre towards a given parameter, the larger improvement that parameter requires. More information about this is also given in Section IV.3.
From this we can see that comparatively little improvement is needed in server efficiency and emission fidelity. The server efficiency for the ion trap considered here is enhanced due to the use of a cavity [48]. Optimising this further is less important in reaching the error probability threshold than increasing the coherence time because we are in a regime where the time most qubits spend in memory is limited by the cutoff, thereby increasing the efficiency will lead to a higher rate (not considered here) rather than a lower error probability.
More improvements are needed for the entangling and single-qubit gate fidelities, and coherence time. As can be seen from Figure 3, many single-qubit gates are executed in order to perform the $CZ$ gate used for creating the graph state, which makes any imperfections in the execution of such single-qubit gates more impactful. We do note that the number of single-qubit gates may be reduced by optimising the graph state creation and perhaps absorbing some of the gates into the measurement bases, which would reduce the overall errors and hence lessen the need for improvement on this parameter. This is not the case for the entangling-gate fidelity.
IV Methods
Here, we discuss some details of how the trapped-ion server and client apparatus are modelled in our NetSquid simulation available at [54]. We also outline the minimisation method used to determine the requirements mentioned in the previous section.
IV.1 Trapped-ion modelling
We model the trapped-ion server using the NetSquid-based library NetSquid-TrappedIons [25]. This library has been first used in [27] and is here introduced in more detail.
We model the decoherence of trapped-ion qubits over time using the collective Gaussian dephasing process described in [55], which can be rewritten as [27]
[…]
with
$K_{r}=\exp\left(-ir\frac{t}{\tau}\sum_{j=1}^{n}\sigma_{z}^{(j)}\right)$ (4)
and
$p(r)=\frac{1}{\sqrt{2\pi}}e^{-r^{2}/2}.$ (5)
[…]
The emission of entangled photons from an ion trap is modelled in NetSquid-TrappedIons as the creation of a photon that has a polarisation degree of freedom that is maximally entangled with the state of the ion used in the emission, |$\ket{\psi}=(\ket{0H}+\ket{1V})/\sqrt{2}$, followed by the application of a single-qubit depolarising channel on the photon’s polarisation. This results in a Werner state of the form $\frac{3}{4}F\ket{\psi}\bra{\psi}+\frac{4F-1}{3}\frac{\mathbb{1}}{4}$, where $F$ is the fidelity of the state to the perfect state $|\psi\rangle$.
IV.2 Client apparatus modelling
The client is modelled as an optical setup where an incoming ion-entangled photonic qubit is transmitted through a half waveplate (HWP) followed by two quarter waveplates (QWPs) to rotate the polarisation state. The photon is then guided to a polarising beam splitter and detected with single-photon detectors.
In the simulation, the effect of the waveplates is given by the multiplication of the waveplate Jones matrices (IV.2) with the state vector [56]. The relative phase retardation induced between the fast axis and the slow axis is $\delta=\pi/2$ for a quarter waveplate and $\delta=\pi$ for a half waveplate. The fast axes waveplates also have an angle $x$ with respect to the x-axis, determining the specific rotation that is implemented. However, errors in the setup can influence the retardation, giving a retardation deviation of $\Delta\delta$. It is also possible to have a deviation in the angle $x$ leading to $\Delta x$. With this, we can write the Jones matrices in general form [57] to include the errors as
$U(\delta^{\prime},x^{\prime})=e^{-i\delta^{\prime}/2}\times$
$\begin{pmatrix}\cos^{2}x^{\prime}+e^{i\delta^{\prime}}\sin^{2}x^{\prime}&(1-e^{i\delta^{\prime}})\cos x^{\prime}\sin x^{\prime}\\ (1-e^{i\delta^{\prime}})\cos x^{\prime}\sin x^{\prime}&\sin^{2}x^{\prime}+e^{i\delta^{\prime}}\cos^{2}x^{\prime}\end{pmatrix},$ (8)
where $\delta^{\prime}=\delta+\Delta\delta$ (with $\delta=\pi/2$ for a QWP, and $\delta=\pi$ for a HWP) and $x^{\prime}=x+\Delta x$. Together, $\Delta\delta$ and $\Delta x$ lead to the waveplate error probability as given in Table 1. The waveplates are implemented as a custom operation in NetSquid according to the Jones matrix.
[30] gives an overview of general qubit rotations in terms of these fast axis settings. We can set the fast axes to correspond to a measurement in the |±$|\pm_{\theta}\rangle$-basis as
$x_{1}=0;$
$x_{2}=\theta/2;$
$x_{3}=\theta/4-3\pi/4.$
IV.3 Minimisation methods
In our analysis, we use one target metric. This target metric is the error tolerance of 25% that has been introduced before. This is a bound on the fraction of test rounds that are allowed to fail while still being $\epsilon$-composably-secure. In [20] the fraction of failed test rounds $w$ over the total number of test rounds $t$ is bound by
$w/t<\frac{1}{k}\frac{2p-1}{2p-2}.$ (9)
Here, $k$ is the principal colouring of the computation graph (which is the smallest number of ‘colours’ or labels one can give to a the nodes in a graph to where no two neighbouring nodes have the same colour, see for example [58]) and $p$ is the inherent error probability of the bounded-error quantum computation. If the error probability is above this fraction, the client aborts the protocol due to security issues. Assuming a computation for which $p=0$ we require the error probability to be below $1/2k$. The one dimensional graph for single-qubit rotations used in this paper is two-colourable, which gives a maximum error tolerance of 25%.
We look for absolute minimal hardware requirements by setting all-but-one parameters to perfect aside from fibre attenuation. For the coherence time, perfect means removing the collective dephasing noise model from the ions in the trap. We then sweep over the non-perfect parameter to find where the error probability due to this imperfection crosses the threshold.
To find the crossing point, we do an initial global search with a small number of test rounds per point (to limit computation time), to find the approximate regime of where the error probability would pass the threshold. Once the regime is located, the search is focused by taking the closest point above and below the threshold, halving the distances to their mean and running the simulation again for these points with a larger number of rounds. This process is repeated until the error-probability confidence intervals of the points crosses the threshold of 25% (this is similar to the bisection method used in root finding). The confidence intervals are determined by Hoeffding’s bound [59] as $\sqrt{\log{(2/0.05)}/2t}$ with $t$ the number of test rounds. The focused search is executed with 70000 points, in order to have a confidence level of 95% in an interval $\pm$ 0.005 for the error probability. The minimal requirements given in Figure 4 are then extracted from the closest points ($(x_{1},y_{1}),(x_{2},y_{2})$) by a linear interpolation at the threshold ($y=0.25$) as $x=(y-y_{1})*(x_{2}-x_{1})/(y_{2}-y_{1})$. The error in these estimates is found by applying the same interpolation to the edges of the error probability confidence interval.
Next, we find the set of minimal improvements. That is, from a given baseline (Table 2), what parameters allow for fulfilling the constraint on the error probability with the least improvement. In order to quantify the cost of improving a parameter by a certain amount, we define hardware costs $H_{c}$ and a cost function $C$, which combines the hardware costs with our constraint on the error probability to give a single-objective minimisation problem as done in [26, 60, 61, 27]. We then employ a genetic algorithm using a workflow manager called YOTSE [62] to find the set of hardware parameters that minimises the cost function. This optimisation was run on SURF’s high- performance-computing cluster Snellius (Platinum 8360Y CPU @ 2.4GHz, maximum of 480 GB RAM) and on a workstation featuring an Intel Xeon Gold 6230 CPU @ 2.10GHz and 188 GB of DDR4 RAM, taking around 100 hours per optimisation run of 20 generations.
To have a consistent way of calculating hardware cost, we associate a probability of no-imperfection $p_{NI}(b_{i})$ to each of the $N$ baseline hardware parameters $b_{i}\in B$, where $B$ is the baseline set of $N$ hardware parameters. This scales all parameters from 0 to 1, where 1 means a perfect setting (e.g., infinite coherence time, 100% efficiency). We can improve upon this baseline with an improvement factor $k$ to find $p_{NI}(x_{i})=\sqrt[k]{p_{NI}(b_{i})}$. These $p_{NI}$ are then summed over for all hardware parameters $x_{i}$ to find the total hardware cost $H_{c}(X)$ of a setup with hardware set $X=\{x_{i}\}_{0<i\leq N}$ with respect to the baseline $B$ as
$H_{c}(X)=\sum_{i=1}^{N}\frac{\ln\{p_{NI}(b_{i})\}}{\ln\{p_{NI}(x_{i})\}}.$ (10)
This is equivalent to summing over the improvement factor of each parameter.
The probabilities of no-imperfection are defined for the optimisation parameters as in Table 3.
Parameter $\mathbf{p_{NI}}$
Server efficiency $\eta$
$\eta$
Coherence time $T_{c}$
$e^{-t^{2}/T_{c}^{2}}$
(Gate and entangled state) fidelity $F$
$\frac{1}{3}(4F-1)$
Table 3: Probabilities of no-imperfection: re-scaling parameters from zero to one. This is used to associate an improvement factor from the baseline to each parameter that is dimensionless and thus comparable.
For derivations and more explanation of the probabilities of no-imperfection, see Supplementary Note 6 of [27]. Note that the $t$ in the probability of no-imperfection of the coherence time, indicating the timescale over which the decoherence happens, does not influence the hardware cost, as $\ln\left(p_{NI}(b_{i})\right)/\ln\left(p_{NI}(x_{i})\right)$ = $\ln\left(e^{-t^{2}/T_{c}^{2}}\right)/\ln\left(\sqrt[k]{e^{-t^{2}/T_{c}^{2}}}\right)$ = $\left(-t^{2}/T_{c}^{2}\right)/\left(-t^{2}/kT_{c}^{2}\right)=k$.
We now combine our requirement of having an error probability below 25% with the hardware cost to find the total cost of a set of parameters. We want the cost assigned to a set of parameters to be very high when the constraint is not met, and to be lower the closer parameter sets are to the baseline assuming that the constraint is met. A function to capture this behaviour, similar to what is used in [26, 27, 60, 61], is
$\begin{split}C=&w_{1}(1+(w/t-1/2k)^{2})\Theta(w/t-1/2k)\\ &+w_{2}Hc(x_{1},...,x_{N}),\end{split}$ (11)
where $\Theta(x)$ is the Heaviside step function and $w_{1}$ and $w_{2}$ are the weights of the objectives. We choose $w_{1}\gg w_{2}$ in order for the function to reflect that it is much more important to satisfy the requirement than to minimise hardware requirements, i.e., we do not care about the hardware cost as long as the requirement is not met.
To minimise the cost function, we use the YOTSE [62] workflow manager, which uses a genetic algorithm. In this, for all parameters, a number of points is drawn from a range between the baseline (Table 2) and their perfect value (except for the coherence time, which is capped at 1 second), i.e., for which $p_{NI}=1$. We initially draw 3 points for server efficiency, 4 for coherence time, 2 for single-qubit fidelity, 3 for entangling gate fidelity and 2 for emission fidelity, meaning the initial population is formed by $3\times 4\times 2\times 3\times 2=144$ sets of parameters. The number of points which are drawn for each parameter is based on the size of the range baseline to perfect for that parameter (e.g., there are more possible values that the entangling gate fidelity can take on than the single-qubit gate fidelity can, as it is further from perfect, we therefore initially draw more points at random from the entangling gate fidelity distribution). From this initial population the lowest-costing 8 ‘parents’ are taken to recombine with a mutation probability of .2 into a new generation. This is repeated over 20 generations. Each point consists of 20000 test rounds for a confidence interval around the error probability of ±$\pm 0.0096$. The minimal costing set of parameters is then fed into a local search algorithm, who decreases the cost of each parameter slightly until the error probability requirement is no longer met. The local search is done with 70000 rounds for a confidence interval of ±$\pm 0.0051$. The outcome of this is a set of parameters that is minimal in the sense that worsening any parameter at that point will result in the requirement not being met.
V Future work
One could modify the cost function (11) to include a constraint on the rate at which the computation or test rounds can be performed. This might change the directions along which the hardware should be improved (i.e., it changes the set of minimal improvements) as parameters such as server efficiency become more important, and conversely the coherence time would become important. This could be done in the same workflow, by changing the cost function and choosing a rate constraint. We have however chosen to not include this, as there is no immediate clear goal in rate, whereas there is a clear goal in error probability (25%).
The constraint on the error probability of 25% we consider in this paper is a theoretical limit. In reality, an error probability of 25% would require an impractically large number of rounds in order to find a desirable $\epsilon$ for security. Instead of setting a constraint on a minimum rate at which the computations can be performed, we might want to find a different metric to target than just the rate and error probability. Having a lower error probability would require one to perform less rounds of the protocol, thereby finishing the total computation faster. An option could be to consider the rate of successfully completed computations, which depends on the error probability as this determines the number of rounds to be performed, as well as on the rate at which these rounds are performed. This is beyond the scope of this paper.
Another thing to consider are requirements for larger graph states. The current analysis considers a universal resource for single-qubit rotations, but a universal resource for any quantum computation (i.e., including two-qubit gates), such as a brickwork state [15], would require more qubits. This will lead to more stringent requirements on the hardware. In principle, the same framework used in this paper could be extended to larger graphs such as the brickwork state, but the current status of the code would make the computation time unreasonably long. The code could be sped up in general by including a framework similar to NetSquid’s entanglement generation Magic [63], which offers simulation speedup through state insertion. An equivalent to this for RSP is currently in development.
The protocol could be optimised by limiting the number of single-qubit gates used in the formation of the graph state, either by considering a full-graph optimisation or by absorbing some of the single-qubit gates into the measurement bases. We also note that, depending on the graph state, not all qubits need to be ‘alive’ in the memory at the same time. The first qubit can be measured before the last qubit is initiated, as long as they are not nearest-neighbours in the graph. In addition, if enough memory qubits are available, the RSP phase could be parallelised by sending ion-entangled photonic qubits from different memory positions successively without waiting for an heralding signal in between.
Data and code availability
The code used to generate, process and plot the data is available at [54], the data is available at [53].
Acknowledgements
We would like to thank Harold Ollivier and Maxime Garnier for useful discussions on rVBQC. We thank Viktor Krutianskii for useful discussions on trapped-ion devices and for sharing experimental data and parameters. We thank David Maier for providing technical support for the code and execution thereof. We thank Kian van der Enden for critical reading of the manuscript. This project (QIA-Phase 1) has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101102140.
Author contributions
JvD led the development and execution of the simulation of the protocol, gathered and analysed the results, and prepared this manuscript. GA contributed to development of hardware models of trapped-ions and prepared Section IV.1 of this manuscript. TzBP, JAS provided feedback and discussions throughout the project. FFdS provided technical support and guidance. TEN helped in the technical understanding of the ion traps and the parameters. All authors revised the manuscript. SW conceived and supervised the project.
Appendix A Formal protocol
Based on [20], as in [22].
Clients inputs: Angles $\{\phi_{v}\}_{v\in V}$ for all qubits (vertices) $V$, determining the gate(s); a graph $G$; a flow $f$ on $G$, determining the order of measurements.
Protocol:
  1. 1.
    The client chooses uniformly at random a partition $(C,T)$ of the set of indices of all the rounds in the protocol $N,C\cap T=\emptyset$, with $C$ and $T$ the set of indices for computation and test rounds, respectively.
  2. 2.
    For all rounds $n\in N$ the client and server perform the following subprotocol (the client may send a message redo n to the server before step 2.3, or the server may send it at any time. Both parties then restart round $n$ with fresh randomness, for more information about this redo feature, see [20]):
    1. 2.1.
      If $n\in T$ (test round), the client chooses uniformly at random a colour $K_{n}$ to define the set of trap vertices for this test round. (The colouring of a graph refers to a way of labelling the nodes such that no neighboring nodes have the same colour, this ensures that no two traps are connected through an edge).
    2. 2.2.
      For all $v\in V$ (i.e., for all qubits): # RSP
      1. 2.2.1.
        The server prepares a bell pair $|\psi\rangle=(|00\rangle+|11\rangle)/\sqrt{2}$ and sends half of it to the client along with a classical ID.
      2. 2.2.2 a.
        When the ID arrives at the client, and the qubit also arrived, the client sends a classical confirmation back to server and performs a measurement yielding outcome $m_{v}$. The measurement basis is chosen as:
        1. (i)
          If $n\in T$ and $v\notin K_{n}$, measure in the standard basis (i.e. prepare a dummy qubit).
        2. (ii)
          If $n\in C$ (computation round) or if $n\in T\land v\in K_{n}$ (trap qubit in test round), measure in ±$\pm_{\theta_{v}}$-basis, with $\theta_{v}\in\{k\pi/4\}_{0\leq k\leq 7}$ randomly chosen, (i.e. prepare |±$|\pm_{\theta_{v}}$).
      3. 2.2.2 b.
        When ID arrives at client, if qubit did not arrive: client sends classical ’lost’ message to server. The server removes the qubit from its memory and goes back to step 2.2.1.
    3. 2.3.
      The client sends description of $G$, the server performs a $CZ$ gate between all qubits that share an edge according to $G$ (i.e. the server constructs the graph state).
    4. 2.4.
      For all $v\in V$ the client and server perform the following subprotocol:
      1. 2.4.1.
        the client instructs the server to measure in a ±$\pm_{\delta_{v}}$ basis defined by:
        1. (i)
          If $n\in C$:
          $\delta_{v}=\phi_{v}^{\prime}+\theta_{v}+m_{v}\pi+r_{v}\pi,$ (12)
          where $r_{v}\in_{R}\{0,1\}$ is chosen uniformly at random. The angle $\phi_{v}^{\prime}$ is defined as
          $\phi_{v}^{\prime}=(-1)^{s_{X,v}}\phi_{v}+s_{Z,v}\pi,$ (13)
          with
          $s_{X,v}=\bigoplus_{l\in S_{X,v}}s_{l},\;s_{Z,v}=\bigoplus_{l\in S_{Z,v}}s_{l}.$ (14)
          Where $\bigoplus_{l\in S_{X(Z),v}}$ represents a modulo 2 summation over the $X$ ($Z$) dependency set for qubit $v$. The dependency sets are defined by $S_{X,v}=f(v-1)$ and $S_{Z,v}=\{l:v\in N_{G}(f(l))\}$, with $N_{G}$ referring to the neighbors of a qubit, which are all other qubits connected to it through an edge. These adaptations to the measurement angles eliminate the need for bit flip and phase corrections in between the measurements. For more background on how the measurement angles are determined ($\phi_{v}$) and adapted ($s_{X,v},s_{Z,v}$), see [23].
        2. (ii)
          If $n\in T\land v\in K_{n}$ (trap):
          $\delta_{v}=\theta_{v}+m_{v}\pi+r_{v}\pi,$ (15)
          i.e. the qubit is being measured in the basis in which it is prepared.
        3. (iii)
          If $n\in T\land v\notin K_{n}$ (dummy):
          $\delta_{v}\in\{k\pi/4\}_{0\leq k\leq 7},$ (16)
          is randomly chosen.
      2. 2.4.2.
        The server measures in the basis defined by the client and and sends back the measurement outcome $b_{v}$.
    5. 2.5.
      For all $\{v:n\in T\land v\in K_{n}\}$ the client verifies that $b_{v}=r_{v}\oplus d_{v}$, where $d_{v}=\bigoplus_{i\in N_{G}(v)}d_{i}$ is the sum over the measurement outcomes of the neighbouring dummies of qubit $v$. If this is false for any trap qubit in the test round, the test round fails. If the number of test rounds exceeds a certain fraction $w/t$, the client aborts. Here, $w/t<\frac{1}{k}\frac{2p-1}{2p-2}$, introduced in (9), is the error threshold to guarantee variability and correctness.
    6. 2.6.
      For all $n\in C$, let $y_{c}$ be the classical output of computation round $c$, the clients checks for a majority output, i.d. checks if there exists a $y$ such that |$\left|\{y_{c}:y_{c}=y\}\right|>\left|C\right|/2$. If there is a majority output, this $y$ is taken as the protocol output and the client sends an OK to the server.
Appendix B Alternative route
Multiple sets of parameters can reach the error probability threshold of 25% for a similar cost. In Section III we give a cost-minimised set of parameters (Figure 5) and discuss what what variations in parameters would yield a similar result. The minimised results show less improvement needed in server efficiency and emission fidelity and more in coherence time, entangling and single qubit gate fidelity. How much improvement is needed in these three main objectives varies slightly over different optimisation outcomes. Here, we provide an alternative optimisation outcome compared to the one in the main text to support this observation in Figure 6. In particular, this minimisation puts slightly more emphasis on the single qubit gate fidelity and a little less on the coherence time and entangling gate fidelity compared to the set given in the main text. In addition, the emission fidelity requires more improvement compared to the solution presented in the main text.
Parameter Baseline Minimally improved value
Server efficiency 0.133 0.594
Single qubit gate fidelity 0.99 0.998
Entangling gate fidelity 0.95 0.986
Emission fidelity 0.974 0.988
Coherence time [ms] 62 103
Figure 6: (a): Alternative set of parameter values required to perform a 5-qubit linear graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km. The requirement for protocol to succeed is having an error probability below 25%. This set minimises the cost function (11), meaning it is closest to the state-of-the-art baseline. A visual representation of these parameters in terms of an improvement factor is given in the bottom figure. (b): Directions along which hardware must be improved to perform a 5-qubit linear graph rVBQC protocol on a trapped-ion server and a measurement-only client over 50 km. The further away the line is from the centre towards a given parameter, the larger improvement that parameter requires. The improvements are given in terms of an improvement factor k, which tends to infinity as a parameter tends to its perfect value and is 1 for no improvement. More information on this can be found in Section IV.3.